# June 13, 2024

### v2

@@ 3.3 Attack Stage: Generate the AE $\mathbf{x}^{\prime}$
Upon obtaining the well-trained $\mathcal{G}$ and $\mathcal{G}^{*}$, in the subsequent attack stage, we can utilize them to generate AEs in a black-box setting. As shown in Fig. 1(b), to perform a black-box attack against the victim model $\mathcal{F}$ with a clean image $\mathbf{x}$, we maintain the VF for decoding the perturbed image $\hat{\mathbf{x}}_{f}$ to be identical to that of the initial image $\hat{\mathbf{x}}$, i.e. $\hat{\mathbf{z}}_{v}$ that is obtained by:
@@
This makes $\hat{\mathbf{x}}_{f}$ remain in close proximity to the clean image as much as possible. Meanwhile, our objective is to identify an AF $\hat{\mathbf{z}}_{a}$ that, when combined with $\hat{\mathbf{z}}_{v}$, has the ability to deceive $\mathcal{F}$. For example, in untargeted attacks, our attack objective targets at finding an image $\hat{\mathbf{x}}_{f}$ by:
$\hat{\mathbf{x}}_{f}=\mathcal{D}^{*}(\mathbf{M}^{*}(\hat{\mathbf{z}}_{v}||\hat{\mathbf{z}}_{a})),$ (14)
$s.t.\ \mathrm{\arg\max_{d}}\mathcal{F}(\hat{\mathbf{x}}_{f},d)\neq y\ \text{and}\ ||\hat{\mathbf{x}}_{f}-\mathbf{x}||_{p}\leq\epsilon.$
Here $y$ denotes the ground-truth label of $\mathbf{x}$. In targeted attacks, the first constraint is transformed to $\mathrm{\arg\max_{d}}\mathcal{F}(\hat{\mathbf{x}}_{f},d)=y$ with $y$ being the target class. Note that in E.q. (LABEL:eq:attackobjpre) and (LABEL:eq:attackobj) we use modules of $\mathcal{G}^{*}$ rather than $\mathcal{G}$, and the reason will be clarified later.
For seeking out such a $\hat{\mathbf{z}}_{a}$ to meet E.q. (LABEL:eq:attackobj), firstly, we have to determine the initial image $\hat{\mathbf{x}}$ for the attack stage so as to acquire $\hat{\mathbf{z}}_{v}$ like in E.q. (LABEL:eq:attackobjpre). Considering the adversarial transferability across different models, we leverage an AE rather than the clean image to initialize a more aggressive AF, expecting to enhance attack efficiency. This initialization strategy is feasible with known class information and available surrogate models in closed-set scenarios, since we can directly generate a transferable AE through a white-box attack method against surrogate models to serve as the initial image. Conversely, in open-set scenarios, due to the uncertainty of class categories, a clean image is more suitable for initializing the attack, i.e. $\hat{\mathbf{x}}=\mathbf{x}$. We will next elaborate on the attack processes separately for both closed-set and open-set scenarios.
In closed-set scenarios, as illustrated in Fig. 1(b), we firstly generate a transferable AE $\hat{\mathbf{x}}$ of $\mathbf{x}$ against an ensemble of local surrogate models using a transferable white-box attack $\mathcal{A}$. In this work, we choose SI-NI-TI-DIM Lin et al (2019) as $\mathcal{A}$, which is one of the most powerful transferable methods. $\mathcal{A}$ could be replaced as other attack methods with stronger transferability in the future. Formally, the generation of the initial image at the untargeted attack stage is expressed as:
@@
$s.t.\ \mathrm{\arg\max_{d}}\mathcal{C}_{j}(\hat{\mathbf{x}},d)\neq y\ \text{and}\ ||\hat{\mathbf{x}}-\mathbf{x}||_{p}\leq\epsilon,\forall\mathcal{C}_{j}\in\mathbb{C}.$
@@
For targeted black-box attacks, this formulation only needs to replace the constraint $\mathrm{\arg\max_{d}}\mathcal{C}_{j}(\hat{\mathbf{x}},d)\neq y$ where $y$ represents the ground-truth label with $\mathrm{\arg\max_{d}}\mathcal{C}_{j}(\hat{\mathbf{x}},d)=y$ where $y$ denotes the target label. Then we can combine E.q. (LABEL:eq:init) and (LABEL:eq:attackobjpre) to acquire $\hat{\mathbf{z}}_{v}$, and maintain the VF for decoding the perturbed candidate image $\hat{\mathbf{x}}_{f}$ to be identical to the VF of $\hat{\mathbf{x}}$, i.e. $\hat{\mathbf{z}}_{v}$. Note that $\hat{\mathbf{z}}_{v}$ in E.q. (LABEL:eq:attackobjpre) is disentangled by $\mathcal{G}^{*}$ but not $\mathcal{G}$ since $\hat{\mathbf{x}}$ belongs to the adversarial domain. Nextly, we have to identify proper AFs $\hat{\mathbf{z}}_{a}$ that, when fused with $\hat{\mathbf{z}}_{v}$, has the ability to deceive $\mathcal{F}$. Therefore, we aim to discover a perturbation $\delta$ whose corresponding perturbed image $\hat{\mathbf{x}}+\delta$ can be used to extract such AFs. Formally, we obtain $\hat{\mathbf{z}}_{a}$ via:
$\hat{\mathbf{z}}=\mathcal{E}^{*}(\hat{\mathbf{x}}+\delta),$ (16)
$\hat{\mathbf{z}}_{v0}=\mathbf{V}_{0}^{*}(\hat{\mathbf{z}}),$
$\hat{\mathbf{z}}_{a0}=\mathbf{A}_{0}^{*}(\hat{\mathbf{z}}),$
$\hat{\mathbf{z}}_{a1v}=\mathbf{A}_{1v}^{*}(\hat{\mathbf{z}}_{v0}),$
$\hat{\mathbf{z}}_{a1a}=\mathbf{A}_{1a}^{*}(\hat{\mathbf{z}}_{a0}),$
$\hat{\mathbf{z}}_{a}=\mathbf{M}_{a}^{*}(\hat{\mathbf{z}}_{a1v}||\hat{\mathbf{z}}_{a1a}).$
Then in close-set scenarios, the perturbed AE $\hat{\mathbf{x}}_{f}$ can be reconstructed as follows:
$\hat{\mathbf{x}}_{f}=\Pi_{\epsilon,\mathbf{x}}\Big{(}\mathcal{D}^{*}\Big{(}\mathbf{M}^{*}(\hat{\mathbf{z}}_{v}||\hat{\mathbf{z}}_{a}))\Big{)}.$ (17)
@@
where $\hat{\mathbf{z}}_{v}$ is obtained by E.q. (LABEL:eq:attackobjpre) and keep invariant while $\hat{\mathbf{z}}_{a}$ is iteratively optimized by using E.q. (LABEL:eq:adv1) and continuously updated $\delta$. The operation $\Pi_{\epsilon,\mathbf{x}}$ represents the projection of the input image into the $l_{p}$ ball bounded by $\epsilon$ and centered at $\mathbf{x}$. This forces the adversarial perturbation to remain within the specified constraint. For simplicity, we recast the right-hand side of Eq.(17) as $\mathbf{T}(\hat{\mathbf{x}}+\delta)$. To find an adversarial $\hat{\mathbf{x}}_{f}$, by combining Eq.(17) and using the transformation of variable approach, we formulate our attack objective in E.q. (LABEL:eq:attackobj) in a unified manner as:
$\mathrm{\min}_{\mu}\mathbb{E}_{\mathcal{N}(\delta|\mu,\sigma^{2})}{L}_{adv}(\mathbf{T}(\hat{\mathbf{x}}+\delta),y,v,\mathcal{F},k),$ (18)
$s.t.\ ||\mathbf{T}(\hat{\mathbf{x}}+\delta)-\mathbf{x}||\leq\epsilon,$
@@
where $\hat{\mathbf{x}}_{f}=\mathbf{T}(\hat{\mathbf{x}}+\delta)$ satisfies $\mathrm{\arg\max_{d}}\mathcal{F}(\hat{\mathbf{x}}_{f},d)\neq y$ for untargeted attacks with $y$ representing the ground truth label of $\mathbf{x}$ or $\mathrm{\arg\max_{d}}\mathcal{F}(\hat{\mathbf{x}}_{f},d)=y$ for targeted attacks with $y$ denoting the target class. We sample $\delta$ from a Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ where the mean $\mu$ is to be optimized and the variance $\sigma$ is fixed to be the optimal value through performing a grid search. For optimizing $\mu$, we adopt the natural evolution strategies method to estimate the gradient of the expectation of ${L}_{adv}$ in Eq.(LABEL:eq:algo3) w.r.t $\mu$. When using stochastic gradient descent with a learning rate $\eta$ and a batch size $\tau$, the update rule for $\mu$ is as follows:
@@
Here, $\gamma_{i}\sim\mathcal{N}(0,I)$, and by setting $\delta_{i}=\mu+\sigma\gamma_{i}$, we obtain $\delta_{i}\sim\mathcal{N}(\mu,\sigma^{2})$.
By obtaining a solution $\mu$ to E.q. (LABEL:eq:algo3) using E.q. (19), we simultaneously achieve the ultimate successful AE $\mathbf{x}^{\prime}$ by calculating:
@@
In open-set scenarios, due to the uncertainty of images’ categories, surrogate models are highly likely to predict outputs inconsistent with that of victim models and consequently generate AEs with weak transferabilities or even with side effects. Thus we set the clean image $\mathbf{x}$ as the initial image $\hat{\mathbf{x}}$, i.e.
$\hat{\mathbf{x}}=\mathbf{x},$ (21)
$\hat{\mathbf{z}}=\mathcal{E}(\hat{\mathbf{x}}),$ (22)
$\hat{\mathbf{z}}_{v0}=\mathbf{V}_{0}(\hat{\mathbf{z}}),$
$\hat{\mathbf{z}}_{a0}=\mathbf{A}_{0}(\hat{\mathbf{z}}),$
$\hat{\mathbf{z}}_{v1v}=\mathbf{V}_{1v}(\hat{\mathbf{z}}_{v0}),$
$\hat{\mathbf{z}}_{v1a}=\mathbf{V}_{1a}(\hat{\mathbf{z}}_{a0}),$
Meanwhile, we still extract AFs $\hat{\mathbf{z}}_{a}$ of the perturbed image $\hat{\mathbf{x}}+\delta$ with $\mathcal{G}^{*}$ as E.q. (LABEL:eq:adv1) to keep fused $\mathbf{x}^{\prime}$ lie on the adversarial domain as much as possible. Similar to close-set scenarios, it is easy to obtain the perturbed image $\hat{\mathbf{x}}_{f}$ in open-set scenarios as in E.q. (17), but obtaining $\hat{\mathbf{z}}_{v}$ in E.q. (17) using E.q. (21) and (LABEL:eq:open1) sequentially, and $\hat{\mathbf{z}}_{a}$ calculated by E.q. (LABEL:eq:adv1). Then we can use the same solution as in the closed set scenario mentioned above to obtain a successful AE $\mathbf{x}^{\prime}$.
Table 2: The attack success rate (ASR %), average number of queries (Avg.Q) of test images in targeted attacks with the target class 864 or untargeted attacks for ImageNet. The best and second best Avg.Q are bold and underlined, respectively.
@@
Figure 2: The visualization of AEs produced by MCGSquare Andriushchenko et al (2020) and DifAttack++ (Ours).
4 Experiments
In this section, we compare our DifAttack++ with SOTA score-based black-box attack methods in both close-set and open-set scenarios. Ablation experiments on the cross domain, the HDF module, and the chosen of initial image in the attack stage are also conducted.
4.1 Experiment Setup
4.1.1 Datasets.
We mainly conduct experiments on the large-scale ImageNet-1k (abbreviated as as ImageNet) Deng et al (2009) dataset in close-set scenarios, on Food101 Bossard et al (2014) and ObjectNet Barbu et al (2019) in open-set scenarios. Food101 consists of 101 food categories, and ObjectNet contains natural images with controls for rotation, background, and viewpoint, which poses more difficult challenges for object recognitions. Images in ImageNet, Food101 and ObjectNet are cropped and resized to $224\times 224$, $224\times 224$ and $336\times 336$ for adapting their assoiciated classifiers. When conducting attacks on ObjectNet in open-set scenarios, we exclude 131 classes that overlap with ImageNet, referring to Barbu et al (2019) for more details about these classes.
4.1.2 Classifiers.
For ImageNet, we collect eight pre-trained classifiers with different architectures, i.e. ResNet18, VGG16, GoogleNet, SqueezeNet, ConvNeXtBase, EfficientNet-B3, Swin-V2-T, and ResNet101 provided by torchvision [^2] https://pytorch.org/vision/stable/index.html. Their Top-1 classification accuracies on the ImageNet test dataset are 69.8%, 73.4%, 69.8%, 58.2%, 84.1%, 82.0%, 82.1%, and 81.9% respectively. These eight classifiers is divided into two groups for experiments. The first group, referred to as the “simple group”, consists of the former four classifiers with relatively lower classification accuracy (below 80%) and relatively simple architectures, which were previously used in our DifAttack work. The second group, denoted as the “complex group”, comprises the subsequent four classifiers not being investigated in DifAttack. These models exhibit higher accuracy (above 80%) and possess diverse and complicated architectures. We select three classifiers from each group as surrogate models, while the remaining one in this group serves as the victim model.
For Food101 and ObjectNet, we use well pretrained OpenAI’s CLIP models provided by Ilharco et al (2021) as victim models to further verify the attack ability of DifAttack++ towards foundation models. They are ViT-B/16 trained on public available image-caption data and ViT-H-14-CLIPA-336 trained on DataComp-1B dataset Gadre et al (2024), which achieve 88.7% and 77.4% zero-shot accuracy on Food101 and ObjectNet, respectively.
4.1.3 Comparative Methods.
We compare our DifAttack++ with SOTA approaches, including DifAttack Liu et al (2024a), MCG Yin et al (2023a), SignH Al-Dujaili and O’Reilly (2020), NES Ilyas et al (2018), $\mathcal{N}$Attack Li et al (2019), SimBA-DCT (shorted as SimBA) Guo et al (2019a), BASES Cai et al (2022), P-RGF Cheng et al (2019), Subspace Guo et al (2019b), GFCS Lord et al (2022) and CGA Feng et al (2022). We evaluate them according to their official source codes. Due to the absence of pre-trained models and code, the experimental results of the CGA method on ImageNet are cited from its published paper. Since MCG Yin et al (2023a) could be combined with different kinds of black-box attack, we choose the best query-efficient combination, MCG+SquareAttack Andriushchenko et al (2020), a.k.a MCGSquare, as a representative. The attack performance of MCG integrated with other attack methods lag far behind MCGSquare. Another work Liu et al (2024b) reports remarkable attack performance in their works, but there is unavailable public models or codes.
4.1.4 Parameters.
Following many previous works, we also set the maximum perturbation on ImageNet, Food101 and ObjectNet to 12/255. The maximum query number is all set to 10,000. For the real-world Imagga API, the maximum query number is limited to 500 due to their query limit. In our DifAttack++, we set $\lambda=1$, $\sigma=0.1$, $\eta=0.01$, $k=5$ or 00 for Eq.(8) and (19).
Table 3: The attack success rate (ASR %), average number of queries (Avg.Q) of test images in targeted attacks with the target class 864 or untargeted attacks for ImageNet. The best and second best Avg.Q are bold and underlined, respectively.
@@ 4.1 Experiment Setup
DifAttack 95.0% 3,962 100.0% 454 62.0% 6,555 99.5% 576 50.0% 7,466 99.0% 675 98.0% 4,625 100.0% 774 76.3% 5,704 99.6% 620
DifAttack++ 100.0% 1,520 100.0% 137 90.5% 3,631 100.0% 225 79.5% 4,672 99.5% 264 100.0% 1,627 100.0% 159 92.5% 2,862 100.0% 196

### v3

@@ 3.2 Attack Stage: Generate the AE $\mathbf{x}^{\prime}$
Upon obtaining the well-trained $\mathcal{G}$ and $\mathcal{G}^{*}$, in the subsequent attack stage, we utilize them to generate AEs in a black-box setting. As shown in Fig. 1(b), to perform a black-box attack against the victim model $\mathcal{F}$ with a clean image $\mathbf{x}$, we maintain the VF for decoding the perturbed image $\hat{\mathbf{x}}_{f}^{i}$ to be identical to that of the initial image $\hat{\mathbf{x}}$, i.e. $\hat{\mathbf{z}}_{v}$ that is obtained by:
@@
This makes $\hat{\mathbf{x}}_{f}^{i}$ remain in close proximity to the clean image as much as possible. Meanwhile, our another objective is to identify an AF $\hat{\mathbf{z}}_{a}^{i}$ that, when combined with $\hat{\mathbf{z}}_{v}$, has the ability to deceive the victim model $\mathcal{F}$. For example, in untargeted attacks, our attack objective targets at finding an image $\hat{\mathbf{x}}_{f}^{i}$ by:
$\hat{\mathbf{x}}_{f}^{i}=\mathcal{D}^{*}(\mathbf{M}^{*}(\hat{\mathbf{z}}_{v}||\hat{\mathbf{z}}_{a}^{i})),$ (14)
$s.t.\ \mathrm{\arg\max_{h}}~{}\mathcal{F}(\hat{\mathbf{x}}_{f}^{i},h)\neq y\ \text{and}\ ||\hat{\mathbf{x}}_{f}^{i}-\mathbf{x}||_{p}\leq\epsilon.$
Here $y$ denotes the ground-truth label of $\mathbf{x}$. In targeted attacks, the first constraint is transformed to $\mathrm{\arg\max_{h}}\mathcal{F}(\hat{\mathbf{x}}_{f}^{i},h)=y$ with $y$ being the target class. Note that in Eqs. (LABEL:eq:attackobjpre) and (LABEL:eq:attackobj) we use modules of $\mathcal{G}^{*}$ rather than $\mathcal{G}$, and the reason will be clarified later.
For seeking out such a $\hat{\mathbf{z}}_{a}^{i}$ to meet Eq. (LABEL:eq:attackobj), we first have to determine the initial image $\hat{\mathbf{x}}$ for the attack stage so as to acquire $\hat{\mathbf{z}}_{v}$; similar to Eq. (LABEL:eq:attackobjpre). Considering the adversarial transferability across different models, we leverage an AE rather than the clean image to initialize a more aggressive AF, expecting to enhance the attack efficiency. This initialization strategy is feasible with known class information and available surrogate models in closed-set scenarios, since we can directly generate a transferable AE through a white-box attack method against surrogate models. However, in open-set scenarios, due to the uncertainty of class categories, a clean image would be more suitable for initializing the attack, i.e. $\hat{\mathbf{x}}=\mathbf{x}$. Below, we elaborate on the attack processes separately for the closed-set and open-set scenarios.
3.2.1 Close-set scenarios
In closed-set scenarios, as illustrated in Fig. 1(b), we first generate a transferable AE $\hat{\mathbf{x}}$ of $\mathbf{x}$ against an ensemble of local surrogate models using a transferable white-box attack method $\mathcal{A}$. In this work, we choose SI-NI-TI-DIM Lin et al (2019) as $\mathcal{A}$, which is currently one of the most effective transferable white-box attack methods. Of course, $\mathcal{A}$ could be replaced by other potentially stronger methods that may emerge in the future. Formally, the generation of the initial image at the untargeted attack stage is expressed as:
@@ 3.2.1 Close-set scenarios
$s.t.\ \mathrm{\arg\max_{h}}~{}\mathcal{C}_{j}(\hat{\mathbf{x}},h)\neq y\ \text{and}\ ||\hat{\mathbf{x}}-\mathbf{x}||_{p}\leq\epsilon,\forall\mathcal{C}_{j}\in\mathbb{C}.$
@@
For targeted black-box attacks, this formulation only needs to replace the constraint $\mathrm{\arg\max_{h}}~{}\mathcal{C}_{j}(\hat{\mathbf{x}},h)\neq y$ where $y$ represents the ground-truth label with $\mathrm{\arg\max_{h}}\mathcal{C}_{j}(\hat{\mathbf{x}},h)=y$ where $y$ denotes the target label. Then we can combine Eqs. (LABEL:eq:init) and (LABEL:eq:attackobjpre) to produce $\hat{\mathbf{z}}_{v}$, and maintain the VF for decoding the perturbed candidate image $\hat{\mathbf{x}}_{f}^{i}$ to be identical to the VF of $\hat{\mathbf{x}}$, i.e. $\hat{\mathbf{z}}_{v}$. Note that $\hat{\mathbf{z}}_{v}$ in Eq. (LABEL:eq:attackobjpre) is disentangled by $\mathcal{G}^{*}$ but not $\mathcal{G}$ since $\hat{\mathbf{x}}$ belongs to the adversarial domain. Then, we also need to identify proper AFs $\hat{\mathbf{z}}_{a}^{i}$ that, when fused with $\hat{\mathbf{z}}_{v}$, has the ability to deceive $\mathcal{F}$. Therefore, we aim to discover a perturbation $\delta_{i}$ whose corresponding perturbed image $\hat{\mathbf{x}}+\delta_{i}$ can be used to extract such AFs. Formally, we obtain $\hat{\mathbf{z}}_{a}^{i}$ via:
$\hat{\mathbf{z}}^{i}=\mathcal{E}^{*}(\hat{\mathbf{x}}+\delta_{i}),$ (16)
$\hat{\mathbf{z}}_{v0}^{i}=\mathbf{V}_{0}^{*}(\hat{\mathbf{z}}^{i}),$
$\hat{\mathbf{z}}_{a0}^{i}=\mathbf{A}_{0}^{*}(\hat{\mathbf{z}}^{i}),$
$\hat{\mathbf{z}}_{a1v}^{i}=\mathbf{A}_{1v}^{*}(\hat{\mathbf{z}}_{v0}^{i}),$
$\hat{\mathbf{z}}_{a1a}^{i}=\mathbf{A}_{1a}^{*}(\hat{\mathbf{z}}_{a0}^{i}),$
$\hat{\mathbf{z}}_{a}^{i}=\mathbf{M}_{a}^{*}(\hat{\mathbf{z}}_{a1v}^{i}||\hat{\mathbf{z}}_{a1a}^{i}).$
Then in close-set scenarios, the perturbed AE $\hat{\mathbf{x}}_{f}^{i}$ can be reconstructed as follows:
$\hat{\mathbf{x}}_{f}^{i}=\Pi_{\epsilon,\mathbf{x}}\Big{(}\mathcal{D}^{*}\Big{(}\mathbf{M}^{*}(\hat{\mathbf{z}}_{v}||\hat{\mathbf{z}}_{a}^{i}))\Big{)}.$ (17)
@@
where $\hat{\mathbf{z}}_{v}$ is obtained by Eq. (LABEL:eq:attackobjpre) and keeps invariant while $\hat{\mathbf{z}}_{a}^{i}$ is iteratively optimized by using Eq. (LABEL:eq:adv1) and continuously updated $\delta_{i}$. Here, the operation $\Pi_{\epsilon,\mathbf{x}}$ represents the projection of the input image onto the $l_{p}$ ball bounded by $\epsilon$ and centered at $\mathbf{x}$. This forces the adversarial perturbation to remain within the specified constraint. For simplicity, we recast the right-hand side of Eq.(17) as $\mathbf{T}(\hat{\mathbf{x}}+\delta_{i})$. To find an adversarial $\hat{\mathbf{x}}_{f}^{i}$, by combining Eq.(17) and using the transformation of variable approach, we formulate our attack objective in Eq. (LABEL:eq:attackobj) in a unified manner as:
$\mathrm{\min}_{\mu}~{}\mathbb{E}_{\mathcal{N}(\delta_{i}|\mu,\sigma^{2})}{L}_{adv}(\mathbf{T}(\hat{\mathbf{x}}+\delta_{i}),y,v,\mathcal{F},k),$ (18)
$s.t.\ ||\mathbf{T}(\hat{\mathbf{x}}+\delta_{i})-\mathbf{x}||\leq\epsilon,$
@@
where $\hat{\mathbf{x}}_{f}^{i}=\mathbf{T}(\hat{\mathbf{x}}+\delta_{i})$ satisfies $\mathrm{\arg\max_{h}}\mathcal{F}(\hat{\mathbf{x}}_{f}^{i},h)\neq y$ for untargeted attacks with $y$ representing the ground-truth label of $\mathbf{x}$ or $\mathrm{\arg\max_{h}}\mathcal{F}(\hat{\mathbf{x}}_{f}^{i},h)=y$ for targeted attacks with $y$ denoting the target class. We sample $\delta_{i}$ from a Gaussian distribution $\mathcal{N}(\mu,\sigma^{2})$ where the mean $\mu$ is to be optimized and the variance $\sigma$ is fixed to be the optimal value through performing a grid search. For optimizing $\mu$, we adopt the natural evolution strategies method to estimate the gradient of the expectation of ${L}_{adv}$ in Eq.(LABEL:eq:algo3) with respect to $\mu$. When using stochastic gradient descent with a learning rate $\eta$ and a batch size $\tau$, the updating rule for $\mu$ is as follows:
@@
where $\gamma_{i}\sim\mathcal{N}(0,I)$, and by setting $\delta_{i}=\mu+\sigma\gamma_{i}$, we obtain $\delta_{i}\sim\mathcal{N}(\mu,\sigma^{2})$. When we obtain a feasible solution for $\mu$ in Eq. (LABEL:eq:algo3) using Eq. (19), we are able to sample at least one successful AE $\mathbf{x}^{\prime}$ by calculating:
@@
Table 2: Results of score-based black-box attacks on ImageNet for the “simple group". Attack success rate (ASR %) and average number of queries (Avg.Q) are used as metrics. For target attacks, the class 864 is selected as the target. The best and second best ASR and Avg.Q results are bold and underlined, respectively.
@@
Table 3: Results of score-based black-box attacks on ImageNet for the “complex group". Attack success rate (ASR %) and average number of queries (Avg.Q) are used as metrics. For target attacks, the class 864 is selected as the target. The best and second best ASR and Avg.Q results are bold and underlined, respectively.
@@
DifAttack 95.0% 3,962 100.0% 454 62.0% 6,555 100.0% 605 50.0% 7,466 99.0% 675 98.0% 4,625 100.0% 662 76.3% 5,652 99.8% 599
DifAttack++ 100.0% 986 100.0% 137 92.0% 3,711 100.0% 225 82.0% 4,285 99.5% 264 100.0% 1,495 100.0% 159 93.5% 2,619 99.9% 196
3.2.2 Open-set scenarios
In open-set scenarios, due to the uncertainty of images’ categories, surrogate models are highly likely to predict outputs inconsistent with that of victim models and consequently generate AEs with rather weak transferabilities. In these scenarios, we hence set the clean image $\mathbf{x}$ as the initial image $\hat{\mathbf{x}}$, i.e.
$\hat{\mathbf{x}}=\mathbf{x},$ (21)
$\hat{\mathbf{z}}=\mathcal{E}(\hat{\mathbf{x}}),$ (22)
$\hat{\mathbf{z}}_{v0}=\mathbf{V}_{0}(\hat{\mathbf{z}}),$
$\hat{\mathbf{z}}_{a0}=\mathbf{A}_{0}(\hat{\mathbf{z}}),$
$\hat{\mathbf{z}}_{v1v}=\mathbf{V}_{1v}(\hat{\mathbf{z}}_{v0}),$
$\hat{\mathbf{z}}_{v1a}=\mathbf{V}_{1a}(\hat{\mathbf{z}}_{a0}),$
Meanwhile, we still extract AFs $\hat{\mathbf{z}}_{a}^{i}$ of the perturbed image $\hat{\mathbf{x}}+\delta_{i}$ with $\mathcal{G}^{*}$ as Eq. (LABEL:eq:adv1) to keep fused $\hat{\mathbf{x}}_{f}^{i}$ lie on the adversarial domain as much as possible. Similar to close-set scenarios, it is easy to obtain the perturbed image $\hat{\mathbf{x}}_{f}^{i}$ in open-set scenarios as in Eq. (17); but obtaining $\hat{\mathbf{z}}_{v}$ in Eq. (17) using Eqs. (21) and (LABEL:eq:open1) sequentially, and acquiring $\hat{\mathbf{z}}_{a}^{i}$ by Eq. (LABEL:eq:adv1). Then we can use the same solution as in the closed-set scenarios mentioned above to obtain a successful AE $\mathbf{x}^{\prime}$.
4 Experiments
In this section, we compare our DifAttack++ with SOTA score-based black-box attack methods in both close-set and open-set scenarios. Ablation studies on the number of disentanglement layers, the CD, the HDF module, and the selected initial image in the attack stage will also be conducted.
4.1 Experiment Setup
4.1.1 Datasets.
We mainly perform experiments on the large-scale ImageNet-1k (abbreviated as ImageNet) Deng et al (2009) dataset in close-set scenarios, on Food101 Bossard et al (2014) and ObjectNet Barbu et al (2019) in open-set scenarios. Specifically, Food101 consists of 101 food categories, and ObjectNet contains natural images with controls for rotation, background, and viewpoint, which poses challenges for object recognition. As a common practice, images in ImageNet, Food101 and ObjectNet are cropped and resized to $224\times 224$, $224\times 224$ and $336\times 336$ for adapting their associated classifiers. When conducting attacks on ObjectNet in open-set scenarios, we exclude 131 classes that overlap with ImageNet. More details about these classes can be found in Barbu et al (2019).
4.1.2 Classifiers.
For ImageNet, we collect eight pre-trained classifiers with different architectures, i.e. ResNet18, VGG16, GoogleNet, SqueezeNet, ConvNeXtBase, EfficientNet-B3, Swin-V2-T, and ResNet101 provided by torchvision [^2] https://pytorch.org/vision/stable/index.html. Their top-1 classification accuracies on the ImageNet validation dataset are 69.8%, 73.4%, 69.8%, 58.2%, 84.1%, 82.0%, 82.1%, and 81.9% respectively. These eight classifiers are divided into two groups for experiments. The first group, referred to as the “simple group”, consists of the former four classifiers with relatively lower classification accuracy (below 80%) and simple architectures, which have also been used in DifAttack. The second group, denoted as the “complex group”, comprises the remaining four classifiers, which have not been investigated in DifAttack. These models exhibit higher accuracy (above 80%), with more diverse and advanced architectures. We select three classifiers from each group to compose the set of surrogate models, while the remaining one in the same group serves as the victim model. In addition, for Food101 and ObjectNet, we use well pretrained OpenAI’s CLIP models provided by Ilharco et al (2021) as victim models to further verify the attack ability of DifAttack++ towards foundation models. These two models are ViT-B/16 trained on publicly available image-caption data and ViT-H-14-CLIPA-336 trained on DataComp-1B dataset Gadre et al (2024), achieving 88.7% and 77.4% zero-shot accuracy on Food101 and ObjectNet, respectively.
4.1.3 Comparative Methods.
We compare our DifAttack++ with SOTA score-based black-box attack approaches, including DifAttack Liu et al (2024b), MCG Yin et al (2023), CGA Feng et al (2022), GFCS Lord et al (2022), BASES Cai et al (2022), SignH Al-Dujaili and O’Reilly (2020), SimBA-DCT (shorted as SimBA) Guo et al (2019a), $\mathcal{N}$Attack Li et al (2019), P-RGF Cheng et al (2019), Subspace Guo et al (2019b), and NES Ilyas et al (2018). We evaluate them according to their official source codes. Due to the absence of pre-trained models and code, the experimental results of the CGA on ImageNet are directly copied from Feng et al (2022). Since MCG Yin et al (2023) could improve the query efficiency of the black-box attack method being combined with, meanwhile SquareAttack is at the top of the BlackboxBench leaderboard Zheng et al (2023, 2024), we choose the best query-efficient combination (see also Yin et al (2023)), MCG+SquareAttack Andriushchenko et al (2020) dubbed as MCGSquare as a representative method to compare with.
4.1.4 Parameters.
Following many previous works, we set the maximum perturbation on ImageNet, Food101 and ObjectNet to 12/255, and the maximum query number to 10,000. For the real-world Imagga API, the maximum query number is limited to 500 due to the practical constraints. Also, in our DifAttack++, we set $\lambda=1$, $\sigma=0.1$, $\eta=0.01$, $k=5$ or 00 in Eqs.(8) and (19).
Figure 2: The visualization of AEs produced by MCGSquare Andriushchenko et al (2020) and DifAttack++.